OpenStudy (anonymous):

\[\sqrt{x} + \sqrt{x + 16 } = 3 \] find the set of solutions

OpenStudy (maheshmeghwal9):

\[\sqrt{a}+\sqrt{b}=c\]\[\implies a+b+2\sqrt{ab}=c^2.\]\[\implies c^2-a-b=2\sqrt{ab}.\]Square again: -

OpenStudy (maheshmeghwal9):

a = x b = x+16 c = 3.

OpenStudy (ghazi):

you'll get complex roots

OpenStudy (anonymous):

i did that way but i was unable

OpenStudy (anonymous):

dont forgot consider resrtictions ... easily u can see \(x\ge 0\) so \(\sqrt{x}+\sqrt{x+16} \ge 4\) and no real roots

OpenStudy (ghazi):

@eliassaab X is not equal to 1/4 it doesn't satisfy the equation

OpenStudy (anonymous):

@mukushla yes u r correct there is no solution.... but can u expl ur method

OpenStudy (ghazi):

roots are complex and it is -16+\[(-16+\pm i 11.66)/ 4\]

OpenStudy (anonymous):

Sorry I read 16 as 6

OpenStudy (anonymous):

There is no solutions.

OpenStudy (maheshmeghwal9):

no solution complex roots will be rejected as x must be greater or equal to 0

OpenStudy (maheshmeghwal9):

@ghazi :)

OpenStudy (maheshmeghwal9):

sqrt function has domain \[x \ge 0.\]

OpenStudy (ghazi):

he hasn't mentioned the condition..that's why i found the complex roots

OpenStudy (anonymous):

got it...00

OpenStudy (maheshmeghwal9):

it does nt depends on mentioning it is a rule for defining functions too:) if u solve an equation:)

OpenStudy (maheshmeghwal9):

if u solve an equation put the value in original equation & see if the function of equation is defined or nt @ghazi :D

OpenStudy (ghazi):

hmm ..agreed .and substituting value makes no sense..i forgot the domain thing..sorry

OpenStudy (maheshmeghwal9):

no worries just practise i also have to do that:)

OpenStudy (ghazi):

what about the domain of complex roots?

OpenStudy (maheshmeghwal9):

complex roots aren't functions then how can they hve dommain???????

OpenStudy (ghazi):

i mean complex functions? when you find it's roots then how you define that function?

OpenStudy (maheshmeghwal9):

sorry i don't have enough knowledge about that but u can take knowledge @ google.com I mean u can search it there:)

OpenStudy (ghazi):

thank you..i had to ask so thought to post it here...and obviously google is a universal solution for me

OpenStudy (maheshmeghwal9):

yeah:) Best of luck :D